Fermi National Laboratory


Accelerator Update

Wednesday May 12
The day shift began with the Tevatron (Tev) studies, Antiproton Source (Pbar) stacking, beam to MiniBoone.

At 9:46 AM it was reported that the M01RAW skid pump motor will need to be rebuilt or replaced. This inhibits beam to MCenter and MTest.

At 11:00 AM problems with the Booster kicker MKS01 lead to the power supply being replaced. This was finished by 12:49 PM.

Large emittances caused stacking to be disabled from 11:55 AM to 1:34 PM

Shot setup begins at 3:41 PM.

M01RAW was repaired by 8:00 PM

Beam to MCenter at 11:15 PM.

Thursday May 13
The Midnight shift began with store #3487 colliding, stacking, and with beam to MiniBoone and MCenter.

Lost the Tev store at 8:00 AM when the Tev ramp dumped. The ramp dump was traced to a battery backup failure at F2.

The Tev was ramping once again at 9:57 AM.

The Tev then went into damper studies.

Beam off to MCenter to permit A/C duct work in MC7.

At 5:20 PM the Tev quenched at BA and BB.

Quench recovery took until 6:21 PM.

Shot setup began at 7:55 PM.

At 9:45 PM a controlled access was made into Booster to do Rad surveys. This time was also used to replace the Booster notcher supply.

At 10:12 PM, Operations established store #3491 with an initial luminosity of 52.32 E30.

Booster is back up at 10:43 PM, we then returned to stacking.

Beam to MCenter at 11:45 PM.

Friday May 14
The Midnight shift began with store #3491 colliding, stacking, and with beam to MiniBoone and MCenter.

Today's Plans
Reverse Proton studies in Pbar, leave the Tev store in until the Saturday owl shift.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/14/2004   email Fermilab